# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

### v1

@@ Abstract
Methods. We compiled 65 isolated (or field) MSP systems (including lone MSPs) that are well astrometrically determined, and calculated their transverse peculiar (or space) velocities $\bm{v}_{\perp}$ and Galactic heights $z$. Assuming that the Galactic-longitude components $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ are statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), we approached the distribution of the $v_{\mathrm{l}}$ components of $\bm{v}_{\mathrm{sys}}$ by the observed $v_{\mathrm{l}}$ sample. Under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption that $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed, we derived the MSP $v_{\mathrm{sys}}$ distribution from the distribution of the $v_{\mathrm{l}}$ component of $\mathrm{v}_{\mathrm{sys}}$. Based on the derived $v_{\mathrm{sys}}$ distribution, we tested the “stable-$v_{\mathrm{l}}$” assumption with dynamical population synthesis (DPS). Additionally, by matching the observed $z$ and the Galactic-latitude components $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$ to the DPS counterparts, we estimated the initial and the current Galaxy-wide scale heights of isolated MSP systems.
Results. We find that the observed $v_{\mathrm{l}}$ sample can be well described by a linear combination of three normal distributions. Accordingly, the MSP $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian components. Our DPS analysis verified the “stable-$v_{\mathrm{l}}$” assumption in the parameter space of this work, and estimated the initial and the current Galaxy-wide scale heights of isolated MSP systems to be about 0.32 kpc and 0.66 kpc, respectively.
[…]
@@ 1 Introduction
Compact stars are normally the dense remnants of stars that have exhausted their nuclear fuel, resulting in objects such as white dwarfs (WDs), neutron stars (NSs), and black holes (BHs). Distinct formation mechanism(s) have been proposed for different types of compact stars. For example, stellar BHs are believed to form mainly through either direct collapse (of massive stars) or a delayed collapse (e.g. Mirabel, 2017). In the latter scenario, a progenitor star’s core, following a supernova explosion, collapses into a proto-NS first, then succumbs to further collapse into a BH due to fallback accretion (e.g. Zhang et al., 2008). As another major form of compact stars, NSs can be born with core-collapse supernovae (CCSNe), while having alternative formation channels such as accretion-induced collapse (AIC; e.g. Tauris et al., 2013) and WD-WD mergers (e.g. Levan et al., 2006). Moreover, NSs encompass various subgroups such as magnetars and millisecond pulsar (MSP) systems, with each NS subgroup having its primary and alternative formation theories (e.g. Ding, 2022).
@@
As distinct formation channels of compact stars may lead to different natal kicks imparted to the compact star, kinematic investigation of compact stars holds a key to discriminating between different formation channels. This investigation can be realized with precise astrometry using space-based telescopes operating at optical/infrared or through very long baseline interferometry (VLBI) observations made at radio, which has been pursued in the study of stellar BHs (e.g. Mirabel et al., 2002; Willems et al., 2005; Atri et al., 2019) and various NS subgroups (e.g. Tendulkar et al., 2013; Gonzalez et al., 2011; Ding et al., 2023, 2024b, 2024c; Disberg et al., 2024b).
@@
In practice, the birth velocity and position of a compact star or a compact star system (i.e., a gravitationally bound system of $\lesssim 3$ stars hosting a compact star) can be derived with the knowledge of 1) the observed 3-dimensional (3D) position, 2) the observed 3D velocity and 3) the age of the system, by integrating the motion backward over time assuming a model of Galactic potential. However, the age of a compact star system is usually not well constrained, therefore requiring additional assumption about the birth site (e.g. thin Galactic disk) of the compact star system (e.g. Fragos et al., 2009; Atri et al., 2019). Furthermore, most compact star systems, especially pulsar systems, do not have the information of radial velocities $v_{\mathrm{r}}$. As a result, the investigation of the systemic recoil velocities $\bm{v}_{\mathrm{sys}}$ (the recoil velocities of lone compact stars or stellar systems hosting compact stars) of pulsar systems has been mainly limited to kinematically young ($\lesssim 10$ Myr old) systems (e.g. Cordes & Chernoff, 1998; Kramer et al., 2003; Lorimer et al., 2006; Verbunt et al., 2017; Igoshev, 2020; Ding et al., 2024c) in order to minimize the dynamic evolution (of the compact star systems) through the Galactic potential.
@@
On the other hand, for any specific subgroup of kinematically old ($\gtrsim 40$ Myr) compact star systems that do not have $v_{\mathrm{r}}$ information, the observed transverse peculiar velocities $\textit{{v}}_{\perp}$ (also referred to as transverse space velocity, i.e., velocity component tangential to the line of sight, with respect to the standard of rest of the neighborhood) is expected to deviate from the $v_{\mathrm{sys}}$ component tangential to the sightline (e.g. Disberg et al., 2024a). Despite the deviation, it has been proposed by Faucher-Giguere & Kaspi (2006) that the Galactic-longitude component $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ is statistically stable over time, which serves as a critical assumption that can link $v_{\mathrm{l}}$ to the $v_{\mathrm{sys}}$ distribution. For this assumption to be validated in the parameter space of interest, dynamical population synthesis (DPS) is required, which simulates the motion of stellar objects through the Galactic potential assuming initial states (i.e., positions and velocities) (e.g. Cordes & Chernoff, 1997).

### v3

# The systemic recoil velocity distribution and the scale height of field millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

@@ Abstract
Methods. We compiled 64 field MSP systems (including 52 binary MSPs and 12 solitary MSPs) that are well astrometrically determined, and calculated their transverse peculiar (or space) velocities $\bm{v}_{\perp}$ and Galactic heights $z$. Assuming that the Galactic-longitude components $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ are statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), we approached the distribution of the $v_{\mathrm{l}}$ components of $\bm{v}_{\mathrm{sys}}$ by the observed $v_{\mathrm{l}}$ sample. Under the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption that $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed, we derived the MSP $v_{\mathrm{sys}}$ distribution from the distribution of the $v_{\mathrm{l}}$ component of $\mathrm{v}_{\mathrm{sys}}$. Based on the derived $v_{\mathrm{sys}}$ distribution, we tested the “stable-$v_{\mathrm{l}}$” assumption with dynamical population synthesis (DPS). Additionally, by matching the observed $z$ and the Galactic-latitude components $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$ to the DPS counterparts, we estimated the initial and the current Galaxy-wide scale heights of field MSP systems.
Results. We find that solitary field MSPs have similar $v_{\mathrm{l}}$ magnitude as binary ones. Additionally, the observed $v_{\mathrm{l}}$ can be well described by a linear combination of three normal distributions. Accordingly, the MSP $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian components. Our DPS analysis verified the “stable-$v_{\mathrm{l}}$” assumption in the parameter space of this work, and estimated the initial and the current Galaxy-wide scale heights of field MSP systems to be about 0.32 kpc and 0.68 kpc, respectively.
Conclusions. According to the MSP $v_{\mathrm{sys}}$ distribution, $\approx 14$% of all the MSPs born in a globular cluster with the nominal 50 $\mathrm{km\penalty 10000\ s^{-1}}$ central escape velocity can be retained. Therefore, the $v_{\mathrm{sys}}$ distribution of field MSP systems may account for the high number of MSPs discovered in globular clusters, which implies that MSPs in star clusters may follow the same formation channel(s) as field MSP systems.
@@ 1 Introduction
Compact stars are normally the dense remnants of stars that have exhausted their nuclear fuel, resulting in objects such as white dwarfs (WDs), neutron stars (NSs), and black holes (BHs). Distinct formation mechanism(s) have been proposed for different types of compact stars. For example, stellar BHs are believed to form mainly through either direct collapse (of massive stars) or a delayed collapse (e.g. Mirabel17). In the latter scenario, a progenitor star’s core, following a supernova explosion, collapses into a proto-NS first, then succumbs to further collapse into a BH due to fallback accretion (e.g. Zhang08). As another major form of compact stars, NSs can be born with core-collapse supernovae (CCSNe), while having alternative formation channels such as accretion-induced collapse (AIC; e.g. Tauris13) and WD-WD mergers (e.g. Levan06). Moreover, NSs encompass various subgroups such as magnetars and millisecond pulsar (MSP) systems, with each NS subgroup having its primary and alternative formation theories (e.g. Ding22).
@@
As distinct formation channels of compact stars may lead to different natal kicks imparted to the compact star, kinematic investigation of compact stars holds a key to discriminating between different formation channels. This investigation can be realized with precise astrometry using space-based telescopes operating at optical/infrared or through very long baseline interferometry (VLBI) observations made at radio, which has been pursued in the study of stellar BHs (e.g. Mirabel02; Willems05; Atri19) and various NS subgroups (e.g. Tendulkar13; Gonzalez11; Ding23; Ding24; Ding24b; Disberg24).
@@
In practice, the birth velocity and position of a compact star or a compact star system (i.e., a gravitationally bound system of $\lesssim 3$ stars hosting a compact star) can be derived with the knowledge of 1) the observed 3-dimensional (3D) position, 2) the observed 3D velocity and 3) the age of the system, by integrating the motion backward over time assuming a model of Galactic potential. However, the age of a compact star system is usually not well constrained, therefore requiring additional assumption about the birth site (e.g. thin Galactic disk) of the compact star system (e.g. Fragos09; Atri19). Furthermore, most compact star systems, especially pulsar systems, do not have the information of line-of-sight radial velocities $v_{\mathrm{r}}$. As a result, the investigation of the systemic recoil velocities $\bm{v}_{\mathrm{sys}}$ (the recoil velocities of lone compact stars or stellar systems hosting compact stars) of pulsar systems has been mainly limited to kinematically young ($\lesssim 10$ Myr old) systems (e.g. Cordes98; Kramer03; Lorimer06; Verbunt17; Igoshev20; Ding24b) in order to minimize the dynamic evolution (of the compact star systems) through the Galactic potential. We note that $v_{\mathrm{r}}$ are distinct from the so-called “Galactic radial velocity” $v_{\mathrm{R}}$ defined as the radial component of motion in cylindrical Galactocentric coordinates. Unlike $v_{\mathrm{r}}$, $v_{\mathrm{R}}$ can be precisely inferred from the transverse velocity of a compact star when the line of sight (to the compact star) is nearly perpendicular to the star’s radial vector in cylindrical Galactocentric coordinates.
@@
On the other hand, for any specific subgroup of kinematically old ($\gtrsim 40$ Myr) compact star systems that do not have $v_{\mathrm{r}}$ information, the observed transverse peculiar velocities $\textit{{v}}_{\perp}$ (also referred to as transverse space velocity, i.e., velocity component tangential to the line of sight, with respect to the standard of rest of the neighborhood) is expected to deviate from the $v_{\mathrm{sys}}$ component tangential to the sightline (e.g. Disberg24a). Despite the deviation, it has been proposed by Faucher-Giguere06 that the Galactic-longitude component $v_{\mathrm{l}}$ of $\bm{v}_{\perp}$ is statistically stable over time, which serves as a critical assumption that can link $v_{\mathrm{l}}$ to the $v_{\mathrm{sys}}$ distribution. For this assumption to be validated in the parameter space of interest, dynamical population synthesis (DPS) is required, which simulates the motion of stellar objects through the Galactic potential assuming initial states (i.e., positions and velocities) (e.g. Cordes97).
Promising supreme timing stability, MSPs have been extensively studied, and are used to test gravitational theories (e.g. Freire12; Zhu19; Freire24) and detect gravitational-wave background at the nano-Hz regime (e.g. Agazie23a; Antoniadis23a; Reardon23; Xu23; Miles25). With spin periods of only $\lesssim\penalty 10000\$60 ms, MSPs are widely believed to have been spun up during the accretion from their respective donor stars. However, as explained in Section 1.3.1 of Ding22, the unexpectedly large number of MSPs retained in globular clusters (e.g. Pfahl02) and possibly also at the Galactic centre (e.g. Abazajian12) suggest very small $v_{\mathrm{sys}}$ of MSPs (e.g. Boodram22), which reinforces exotic MSP formation channels such as the AIC channel (e.g. Gautam22). Accordingly, it is likely that the MSP $v_{\mathrm{sys}}$ distribution is multi-modal and/or multi-component (see Section 3 or Appendix B.1 for explanation), which has not been indicated by previous studies of MSP kinematics likely due to the relatively large astrometric uncertainties (Hobbs05; Gonzalez11).
Up till now, a well defined MSP $v_{\mathrm{sys}}$ distribution remains largely absent, mainly due to the aforementioned technical challenges (i.e., long-term dynamic evolution and the lack of $v_{\mathrm{r}}$ information) and relatively large astrometric uncertainties. This absence hinders the studies of MSP formation channels, and results in weak constraints on the NS retention fractions in star clusters. Recent years have seen a major increase in the number of well astrometrically determined MSPs (e.g. Ding23; Shamohammadi24), which promises a refined characterization of the MSP $v_{\mathrm{sys}}$ distribution. In this paper, we determine the MSP $v_{\mathrm{sys}}$ distribution with the observed $v_{\mathrm{l}}$ calculated for 64 well astrometrically determined MSPs, then discuss its implications for the NS retention fractions in star clusters. Throughout this paper, uncertainties are stated at 68% confidence.
The establishment of the observed $\bm{v}_{\perp}$ distribution requires precise determination of both proper motions and distances for the MSPs, especially as MSP $v_{\perp}$ are much lower than those of normal pulsars (Hobbs05; Gonzalez11; Matthews16; Shamohammadi24). Applying low-precision astrometric results to MSP kinematics studies would not only shift the average $v_{\perp}$ to the higher end, but also reduce the ability to resolve the potential multiple modes in the $\bm{v}_{\perp}$ distribution. Compared to other methods, the determination of trigonometric parallaxes through VLBI astrometry (e.g. Ding23), pulsar timing (e.g. Perera19; Shamohammadi24), or Gaia astrometry (e.g. Jennings18; Moran23), offers the most precise and reliable distances for MSPs. Recent releases of the astrometric results by Ding23 and Shamohammadi24 have significantly enriched the sample of well astrometrically determined MSPs, which offers a rare opportunity to make statistically significant investigation of the MSP kinematics based purely on well astrometrically determined MSPs.
We searched for viable MSPs within the PSRCAT Catalogue [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester05). Due to relatively frequent dynamic interactions inside globular clusters, the $\bm{v}_{\perp}$ of globular cluster MSPs are likely not reflective of their $v_{\mathrm{sys}}$. With globular cluster MSPs excluded from the study, we reached 64 field MSP systems having significant ($3\,\sigma$) proper motion and parallax determinations (see Table 6). When only focusing on field MSPs with $5\,\sigma$ and $7\,\sigma$ parallaxes, we obtained two high-precision samples of 42 and 30 MSPs, respectively (see Appendix A.1). The details of the sample selection are described in Appendix A.
Following the methods detailed in Section 6 of Ding23, we derived the distances $d$ and the $\bm{v}_{\perp}$ of the 64 MSPs from their parallaxes and proper motions. The $\bm{v}_{\perp}$ have been corrected for the Solar motion and converted to the Local Standard of Rest of each pulsar. In deriving $d$ from the parallaxes, the Lutz-Kelker bias (Lutz73) has been systematically corrected. This correction is particularly important for MSPs with relatively low parallax significances ($\lesssim 6\,\sigma$), as directly inverting parallaxes tends to underestimate their $d$ (and likely $v_{\perp}$ as well). From the $d$, we calculated the Galactic heights $z\equiv d\sin{b}$ for the 64 MSPs, where $b$ refer to Galactic latitudes. The derived $d$, $\bm{v}_{\perp}$ and $z$ are summarized in Table 6, where the columns related to $\bm{v}_{\perp}$ include the $\bm{v}_{\perp}$ magnitudes $v_{\perp}$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of $\bm{v}_{\perp}$. On top of $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, $z$ serves as an additional indicator of MSP kinematics, as the average level of $|z|$ expected to grow with higher $v_{\mathrm{sys}}$. Therefore, this work includes the studies of three samples of measurements: the $v_{\mathrm{l}}$, the $v_{\mathrm{b}}$, and the $z$ samples.
Based on the calculations of the Pearson coefficients, no linear correlation is found between any two of the three samples. Accordingly, the three samples are treated independently. We conducted Wilcoxon signed-rank test on each of the three samples of 64 measurements, and confirmed no indication of asymmetry about 0 in any sample. Accordingly, in the following analysis, we assumed that the samples of the observed $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ are symmetric about 0.
To 1) visualize the three samples of measurements, and 2) investigate the underlying probability density function (PDF) for each sample, we assumed each measurement follows a split normal distribution, simulated 10,000 values per measurement, and concatenated the simulations into a chain of 640,000 values for each of $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$. Despite the simulated sample of 640,000 following essentially a multi-modal distribution, the visualization is expected to characterize where the measurements overlap. The resultant histograms are displayed in Figure 2. According to Table 6 and Figure 2, most of the 64 MSPs are within the scale height of the Galactic thick disk (i.e., $\sim 1$ kpc).
Figure 2: The upper-left, lower-left and lower-right panels display the histograms of Galactic heights $z$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of transverse peculiar velocities, respectively. Each histogram is concatenated from 10000 simulations drawn from the assumed split normal distributions for the 64 measurements (of $z$, $v_{\mathrm{l}}$ or $v_{\mathrm{b}}$; see Table 6). The gray dashed curves show the normalized count smoothed with kernel density estimation (using scipy.stats.Gaussian_kde). The best-fit Cauchy, Laplace and three-component normal distributions are plotted in blue and green dashed curves. Overlaid are the dynamical population synthesis results (labeled as “evolved”) that best match the observed $z$, $v_{\mathrm{l}}$ and $v_{\mathrm{b}}$, based on the determined initial MSP scale height $\zeta_{0}=0.32$ kpc. The corresponding initial distributions are also provided except for $v_{\mathrm{l}}$, as the initial $v_{\mathrm{l}}$ distribution adopted in the dynamical population synthesis is identical to the best-fit three-component normal distribution (a linear combination of three normal distributions).
2.1 Binary vs. Solitary
While most field MSPs are believed to be formed in binary systems where the NSs are spun up during the accretion from their main-sequence companions (e.g. Alpar82; Bhattacharya91), $\gtrsim 20\%$ of field MSPs are found to have no companions. It is proposed that at least some of these solitary (or single) MSPs have ablated their companions after being spun up through accretion (e.g. Fruchter88; Kluzniak88). This scenario is supported by the discoveries of planets around the MSP PSR J1300++1240 (or PSR B1257++12) (Wolszczan92; Konacki03), and planetary-mass companions around the MSPs PSRs J1719-1438 and J2322-2650 (Bailes11; Spiewak18). Additionally, the persistent achromatic component in the timing residuals of PSR J1939++2134 (or PSR B1937++21) has been attributed to the presence of an asteroid belt surrounding the MSP (Shannon13). Of the four aforementioned MSPs supporting the “companion ablation” scenario, PSRs J1300++1240 and J1939++2134 are included in this work, and are classified as solitary MSPs (see Table 6).
In addition to the companion ablation scenario, alternative theories have been proposed for the formation of solitary field MSPs, including (but not limited to) the AIC channel (Tauris13; Freire13) and the strange star scenario (Jiang20a). Particularly, the AIC channel is suggested to result in relatively small peculiar velocities (e.g. Gautam22). Therefore, precise astrometry of solitary MSPs can be used to potentially distinguish between the companion ablation and the AIC channels. Previous astrometric investigations by Hobbs05; Gonzalez11; Matthews16; Shamohammadi24 suggest no indication that solitary field MSPs and binary ones follow different $v_{\perp}$ distributions. Here, we revisit the investigation in better detail, incorporating refined astrometric data and improved analytical techniques.
Among the 64 selected MSPs, 12 are solitary MSPs and 52 MSPs are in binary systems (see Table 6). Following the aforementioned way of visualization, we display the normalized histograms of $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ for the 52 binary MSPs and 12 the solitary MSPs in Figure 4. The global magnitudes of $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ of the solitary and the binary MSPs are summarized in Table 2. We found that the median and dispersion of $|v_{\mathrm{l}}|$ of solitary MSPs well agree with the counterparts of binary MSPs. Assuming 1) the $v_{\mathrm{l}}$ distribution is stable over time (see Section 4 for justification) and 2) MSP systemic recoil velocities are equally possible in all directions, the agreement in $|v_{\mathrm{l}}|$ suggests that solitary field MSPs may follow a systemic recoil velocity distribution similar to that of binary field MSPs. Accordingly, most solitary field MSPs are expected to follow the primary formation channel of binary MSPs (in which they are spun up through accretion from their companions), except that solitary field MSPs eventually evaporate their companions through irradiation. Though the AIC channel is suggested to produce the bulk of the slow-moving MSPs that are presumably retained in the Galactic centre and contribute to the gamma-ray excess observed in the direction of the Galactic centre (Boodram22; Gautam22), the observed $v_{\mathrm{l}}$ suggest that the AIC channel should contribute little to the production of solitary field MSPs. In the following discussions, we assume that solitary and binary field MSPs share the same systemic recoil velocity distribution, and derive the systemic recoil velocity distribution with the full sample of 64 MSPs.
On the other hand, both $|z|$ and $|v_{\mathrm{b}}|$ of solitary MSPs appear to be overall smaller than the counterparts of the binary MSPs. In other words, binary MSPs are found to be more spread out from the Galactic plane (than solitary MSPs), with their $|z|$ nearly following a uniform distribution between 0–1 kpc (see Figure 4). This spread may reflect the observational bias that searching for binary MSPs in the scattered low-Galactic-latitude sky regions is disproportionately harder (compared to solitary MSPs), which could be gradually resolved with increasingly advanced binary pulsar search techniques (e.g. Balakrishnan22; Sengar25). As the sample of solitary field MSPs remains small, we use the full sample of 64 MSPs to estimate the initial and the evolved scale heights of field MSPs in Section 4, which should be considered the weighted average values of the two sub-samples (of solitary and binary MSPs).
Figure 4: The curves in the upper-left, lower-left and lower-right panels display the normalized histograms of the observed Galactic heights $z$, Galactic-longitude component $v_{\mathrm{l}}$ and Galactic-latitude component $v_{\mathrm{b}}$ of transverse peculiar velocities, respectively, for 52 binary MSPs and 12 solitary MSPs (see Table 6). The histograms have been smoothed with kernel density estimation using scipy.stats.Gaussian_kde. Each histogram is concatenated from 10000 simulations drawn from the assumed split normal distributions for the measurements (of $z$, $v_{\mathrm{l}}$ or $v_{\mathrm{b}}$; see Table 6).
Table 2: Global magnitudes of the Galactic heights and the transverse peculiar velocities of the binary and the solitary millisecond pulsars. [^2] The lower bound, the estimate, and the upper bound corresponds, respectively, to the 16th, the 50th, and the 84th percentile of each of the $|v_{\mathrm{l}}|$, $|v_{\mathrm{b}}|$, and $|z|$ chains.
MSPs $|v_{\mathrm{l}}|$ $|v_{\mathrm{b}}|$ $|z|$
($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$) (kpc)
binary $44^{+85}_{-31}$ $34^{+56}_{-24}$ $0.46^{+0.51}_{-0.30}$
solitary $40^{+69}_{-26}$ $17^{+26}_{-13}$ $0.13^{+0.33}_{-0.08}$
2.2 Modal analysis based on different parallax significance
Following the method detailed in Section 7.2.3 of Ding24, we ran Hartigan’s dip tests (Hartigan85) on the three samples (of 64 measurements), and found no indication of multi-modality in any of the samples. As the $v_{\mathrm{l}}$ distribution is believed to be stable over time (Faucher-Giguere06), here we mainly discuss the dip tests on the $v_{\mathrm{l}}$ sample. When assuming that the $v_{\mathrm{l}}$ sample follows a distribution symmetric about 0 (as is supported by the aforementioned Wilcoxon signed-rank tests), the $|v_{\mathrm{l}}|$ sample should, in principle, enhance the indicators of potential multi-modality. Should the $v_{\mathrm{l}}$ sample truly follow the multi-modal distribution used in the visualization, about 180 and 240 additional $v_{\mathrm{l}}$ measurements would be needed to rule out uni-modality of the $|v_{\mathrm{l}}|$ sample at 90% and 95% confidence, respectively, according to our simulations.
Additionally, we ran the same analysis on high-precision MSP samples obtained with higher thresholds of parallax significance $\varpi/\sigma_{\varpi}$ (see Appendix A.1). The results for $v_{\mathrm{l}}$ and $|v_{\mathrm{l}}|$ are summarized in Table 4. According to Table 4, raising $\varpi/\sigma_{\varpi}$ thresholds reduces MSP sample sizes and lowers the level of unimodality of $|v_{\mathrm{l}}|$ or $v_{\mathrm{l}}$ measurements. The latter trend is likely caused by small sample effects. However, it is not unlikely that more high-precision astrometric measurements made with VLBI or pulsar timing could resolve the $v_{\mathrm{l}}$ distribution, and reveal multi-modality of MSP $|v_{\mathrm{l}}|$ with certainty in future. In the following discussions, we only focus on the full MSP sample (of 64), and assume that the $v_{\mathrm{l}}$ distribution is unimodal (given no indication of multi-modality).